Crew 265 Operations Report 24 APR 2022

[Crew 265 Operations Report April 24th]

Crew 265 Operations Report 24-04-2022

SOL: 0

Name of person filing report: Dave Laude

Non-nominal systems: Heating system.

Notes on non-nominal systems: The heating system is now functional after non-functioning with previous crew. I used the first new thermostat that was ordered as that is correct for this system. However, it does not have fan switches like the original did. The main issue is that the exhaust gas fan to the outside cycles on and off because the fan operation pressure sensing switch is stuck closed and so the furnace senses this as a fault. For operation I have to fake the opening and closing during start-up as it is being checked by the electronics. We only use the furnace in the morning for a few hours if very cold in the hab and when I can tend to it. Warm temps are coming soon and so we may not need it. It will not be run overnight as is. No replacement part is planned to be ordered because the entire furnace is to be replaced after this crew season is over.


Spirit rover used: No

Hours: (before EVA)

Beginning charge: (Before EVA)

Ending charge: (On return from EVA, before recharging)

Currently charging: Yes

Opportunity rover used: No


Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: No


Beginning charge:

Ending charge:

Currently charging: Yes

Perseverance rover used: No


Beginning charge:

Ending charge:

Currently charging: Yes

General notes and comments:

Summary of Hab operations:


Water (static tank): Not full. ~480 gallons. Does this warrant a special trip for water

Water (loft tank): 20 gallons

Water Meter: 0158026.1 units

Static to Loft Pump used – No

Static tank pipe heater (on or off): Off

Static tank heater (On or off): Off

Toilet tank emptied: No

Summary of internet: Awesome!

Summary of suits and radios: Batteries charging. One radio in RAM seems to have a very weak battery battery. Will need to test it. Radio replaced with one from lower hab deck. Testing of some currently unused EVA suit battery capacity in inventory has started. Scott Davis has asked for this on a sample basis to confirm or not that batteries labeled dead are of no use. Will start some radio battery capacity testing soon.

Summary of GreenHab operations: Not used, nothing to report

WATER USE: gallons

Heater: On or Off

Supplemental light: On or Off


Summary of ScienceDome operations: Not used, nothing to report

Dual split: Heat or AC, Off

Summary of RAM operations: Tools used for furnace work. DMMs used for battery checks.

Summary of any observatory issues: Won’t be using, nothing to report

Summary of health and safety issues: Star Link cable has been taped to the floor on the lower deck to prevent tripping upon it.

Questions, concerns and requests to Mission Support: (note trash pickup)

Water (static tank): Not full. ~85% or ~470 gallons. Maybe needs another 6-8". Does this warrant a special trip for water tomorrow or can someone deliver some during the mission?

Minutes ago we smelled natural gas or propane on the lower deck only. The gas detector alarm was not set off. Manual detectors read null and the gas dissipated in about 10 minutes. This same situation has occurred on other crews. We will check periodically for the same occurance.

Copyright © The Mars Society. All rights reserved. | Main Site